OpenStudy (anonymous):

help. simplify. the square root of 2/ the square root of 30 - the square root of 2

OpenStudy (alekos):

\[(\sqrt{2}\div \sqrt{30}) - \sqrt{2} \] Is this what you mean?

OpenStudy (alekos):

or this \[\sqrt{2}/(\sqrt{30}-\sqrt{2})\]